If a cost-minimizing firm sells its services in an oligopolistic market, with a marginal product of capital of 10, a rental rate of capital of $5, and a wage rate of $20, what is the marginal product of labor?


4
20
40
60
80
1
Expert's answer
2021-08-06T05:10:35-0400

Solution:

The marginal product of labor = 40


We know that, "\frac{MP_{L} }{MP_{K}}"


Where: w = wage rate

             r = Rental rate of capital

            MPL = Marginal Product of Labor

            MPK = Marginal Product of Capital


"\frac{MP_{L} }{20} =\frac{10 }{5}"


"\frac{MP_{L} }{20} =2"


MPL = 40